12 degrees less than one half the supplement if an angle is the same as the complement. Find the measure of the angle.
Answer by nyc_function(2733)

(180 - A) = supplent of the angle


(90 - A) = complement of the angle


(1/2)(180 - A) - 12 = (90 - A)


Can you finish?